Solve the integral of: x/1 - x^(2) dx
1
Expert's answer
2021-01-18T13:28:23-0500

Let's solve this problem

"\int\large\frac{x}{1-x^2}dx"

Substituting "u = 1-x^2 \to \large\frac{du}{dx} = -2x(steps) \to dx=-\large\frac{1}{2x}du:"

"= -\frac{1}{2}\int\large\frac{1}{u}du"

Now solving: "\int \large\frac{1}{u}du = ln|u|+C"

Plug in solved integrals: "-\frac{1}{2}\int\large\frac{1}{u}du = -\large\frac{ln|u|}{2}+C"

Substituting back "u = 1-x^2:" "\int\large\frac{x}{1-x^2}dx = -\large\frac{ln(1-x^2)}{2}+C"

Answer: "-\large\frac{ln(|x^2-1|)}{2} + C"